OpenStudy (anonymous):

x2+12x−64=0 need help with this problem, so please help

OpenStudy (ashish56):

is it \[x^{2}\] or 2x

OpenStudy (anonymous):

4 , -16

OpenStudy (anonymous):

if it is x^2

OpenStudy (ashish56):

if it is \[x^{2}\] then here is a solution: \[x ^{2}+16x-4x-64 = 0\] \[x(x+16)-4(x+16)=0\] (x+16) (x-4) = 0 x = 4 or x= -16